For the frame pictured below, determine the pin reactions at A, B, and C. 20-0 A 12-0 4 kips E 12-0 10-0 10-0 6 kip
Answer #1

A 121 B Jo 4 kips E 6kps F.B.D of CD B Rey Roy < RBK B RCA c 4 kips equilibrium at oqui & Fx=o. Rexal Roxa Rea 4 Fч со Rey 4мроо Веч о - 4к | ого Pче акiрѕ Reyt Roy 4. Roy = 2 kips FBD of E.CA RAY ARAX. A Qo RLT РЧ ао 6hes & Fyzo о - 6 - Rey + R& Ex=0 RAX Х Rcx=0 RAXE PCX Ž MA=0 6x(12 + lotto) + Rcyx (lotlo) - Reax (20) 6a112+10+10)+2a (10+10) - RLX X 2020 Rcx=11.6KIP2 RB=) Retro = Jili6²+22 Ro=11771 Kips Reaction at c. Rcxa llor Kips. Rey=a kips. Rc=J Reatrey Rc = Jl162 +22 11.77lkips Realplz comment if you have any doubts